# thanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ Abstract
Terahertz radiation finds an increasing number of applications, yet efficient generation and detection remain a challenge and an active area of research. In particular, the precise detection of weak and narrowband terahertz signals is notoriously difficult. Here, we employ a novel type of single-photon detector based on Rydberg atoms to both detect and calibrate a terahertz frequency comb over an octave-spanning range, yet with a MHz-level selectivity. We calibrate the intensity of the electric field of the comb against the fundamental atomic properties, while achieving the intensity (power) sensitivity down to $35.9\text{\,}$ ($1.46\text{\,}$) within a single mode of the frequency comb, all in a room-temperature operated setup. Our results elucidate the transition of terahertz frequency combs into the quantum regime, enabling high-precision and high-sensitivity spectroscopy. This breakthrough allows terahertz science to better leverage revolutionary techniques developed for optical frequency combs.
@@ II.1 Detection
The first, auxiliary method relies upon a 87Rb vapor cell and two optical fields. The laser beams are introduced into the cell from opposing directions and overlaid colinearly. This setup is the typical two-photon excitation scheme utilized in Rydberg atom electrometry, where usually two lasers are used to excite the atoms. The $780\text{\,}$ probe beam tuned to the D2 line is scanned in frequency and after interaction with atomic vapors is shun upon an avalanche photodiode (Thorlabs APD430A) to obtain the probe transmission spectrum. By the observation of the EIT (electromagnetically induced transparency) effect, introduced by the counter-propagating $483\text{\,}$ coupling laser tuned to the $5^{2}\text{P}_{3/2}(\text{F}=3)\rightarrow 27^{2}\text{D}_{5/2}$ transition, the setup is capable of performing RF electric field measurements by analyzing the Autler-Townes splitting in the domain of the probe laser detuning [35].
Figure 1: a) Simplified scheme of the setup used for the calibration and detection of THz frequency comb signal. $\mathsf{TRA}$ – automotive radar chip, $\mathsf{HWP}$ – half-waveplate, $\mathsf{QWP}$ – quarter-waveplate, $\mathsf{PD}$ – photodiode, $\mathsf{PM}$ – parabolic mirror, $\mathsf{PCA}$ – photoconductive antenna, $\mathsf{POL}$ – polarizer, $\mathsf{DM}$ – dichroic mirror, $\mathsf{SF}$ – spectral filter. The removable metallic mirror allows switching between signal sources. The TRA beam has slanted polarizers to discard reflections, while the HWP + POL setup has two copies to increase the achievable attenuation. b) State-level diagram of fields and state transitions facilitating the conversion of frequencies centered around various detected frequencies to the 776 nm output optical signal. The targeted THz signal frequencies are $0.125\text{\,}$, $0.160\text{\,}$, $0.182\text{\,}$ and $0.242\text{\,}$. Note that only one THz frequency is targeted at a time – the coupling and decoupling lasers are coarsely detuned to address the relevant transition. Furthermore, by fine-detuning the decoupling laser by $\Delta_{1200}$ it is possible to tune to the frequencies near the chosen transition.
This measurement scheme has a fundamental limitation on the minimal resolvable field, resulting from the finite EIT linewidth, caused by the movement of atoms in room temperature and the arising Doppler broadening. With the objective of measuring much weaker fields, the A-T setup is then extended by introducing a third laser beam, namely the 1278 nm decoupling field tuned to the $26^{2}\text{F}_{7/2}\rightarrow 5^{2}\text{D}_{5/2}(\text{F}=4)$ transition, realizing the full setup, visible in Fig. 1 a).
@@
Such a detection scheme achieves superior detection sensitivity to the intensity of THz fields, down to the thermal noise limit, with the potential for time-resolved measurements and synchronization with a pulsed source. Thanks to having a strong reference field source, it is possible to calibrate the photon count readout to the electric field amplitude measured from A-T splitting. What is more, as is the case with A-T splitting, we are not limited to resonant RF detection and addressing different frequencies is possible by detuning the coupling or decoupling fields. For simplicity, this works explores tuning just the decoupling laser for mapping the bandwidth of a particular RF transition.
@@ II.2 Signal generation
The THz field is introduced from a high-power source, namely an automotive radar chip TRA_120_045 from Indie Semiconductor. While A-T measurements may be performed with the signal entering the vapor cell from the side, the parametric conversion necessitates phase-matching, so the THz field is combined colinearly with optical fields and focused using a parabolic mirror with a through hole for optical beams. This enables colinear propagation of all fields in the vapor cell, and the setup is further improved using custom-made metamaterials, which allow for beam collimation, attenuation, and polarization control [33]. Both lenses and waveplates are made using an FDM (fused deposition modeling) 3D printer and HIPS (high impact polystyrene) filament, while the polarizers are custom-made flex PCBs (printed circuit boards). All of the designs are fine-tuned for the operating transition frequency using FDTD (finite difference time domain) simulations.
As such, a sufficiently powerful source, like the aforementioned radar chip is employed for calibration and connecting the results obtained in both detection methods. Subsequently, we can switch to conversion detection and study lower intensity field sources, in particular, a THz frequency comb. We therefore introduce a PCA (photoconductive antenna – BATOP iPCA-21-05-300-800-h), which, given a bias voltage, emits THz pulses following the intensity characteristic of a femtosecond Spectra-Physics MaiTai laser, which has a repetition rate $f_{\text{rep}}$ of $79.49\text{\,}$. The PCA has a hyperhemispherical silicon lens mounted on top, which partially focuses the output. It is then collimated and sent into the setup with the same path, and a metal sheet is used as a mirror to toggle between the high-power chip and the locally weaker PCA.
@@ III.1.1 A-T splitting calibration
We begin the calibration by measuring the A-T splitting in the probe transmission spectrum. For this measurement, the decoupling laser is covered, and only the effects from the coupling laser and THz field can be seen. The probe transmission spectrum is then measured for attenuations of the THz field between 0 and $1.5\text{\,}$ introduced by two pairs of HWPs (half-waveplates) and a polarizer. The A-T splitting, created by the THz field, can then be directly observed in the probe transmission spectra as the splitting of the EIT caused by the $483\text{\,}$ laser. Next, we determine the Rabi frequencies for the corresponding splitting by fitting the numerical solutions of the steady state to the measured probe transmission spectra [37].
Subsequently, knowing the dipole moment $d$ of the considered THz transition, the strength of the oncoming field can be expressed in the amplitude of the electric field $E$ by relation
$E=\frac{h}{|d|}\Omega.$ (1)
In the experiment, the dipole moment of the main considered atomic transition, $27^{2}\text{D}_{5/2}\rightarrow 26^{2}\text{F}_{7/2}$, is equal to $d=594a_{0}e$ [38], where $a_{0}$ is the Bohr radius and $e$ is value of the elemental charge. The calibration of the A-T splitting regime is then determined by fitting a linear function with a fixed slope, equal to 1-1, to the calculated electric field amplitudes as a function of THz field attenuation.
Figure 2: Calibration curve derived for the $27^{2}\text{D}_{5/2}\rightarrow 26^{2}\text{F}_{7/2}$ THz transition from the single photons and A-T splitting measurements, together with the fitted calibration line and both the comb tooth signal and the thermal noise levels. Colored areas correspond to the measurement series for the TRA source intensities $I_{1}$ (green) and $I_{2}$ (blue). Respective horizontal colored axes represent attenuation values corresponding to the intensity, whereas the vertical axes represent the photon counts per second measured in both cases. Note that in the $I_{1}$ measurement series a ND 2 filter is inserted at the SPAPD input, which results in factor of 100 lower registered photon counts.
@@ III.1.2 Photon counts calibration
To perform the measurements of converted photon counts, the probe laser is locked at the transition resonant frequency, the decoupling laser is turned on, and both the coupling and decoupling lasers are tuned to maximize the conversion signal coming from the comb tooth closest to the resonance. Then, the photon count measurements of the converted signal are performed for the attenuation range spanning from 0 to 32 dB via the setup of HWPs and polarizers.
@@
To further extend the calibration range, the measurements are done for two TRA source intensities $I_{\text{src}}$: $I_{1}$ used for the A-T splitting measurements, and $I_{2}$, for which the A-T splitting cannot be measured in the probe transmission spectrum. As the conversion signal due to the $I_{1}$ intensity saturates the detector, for those measurement series an additional ND 2 filter is added before the detector. Finally, we turn off the TRA chip and measure the photon counts corresponding to the frequency comb tooth closest to the atomic resonance. Additionally, we cover both THz sources and measure the conversion signal corresponding to the thermal radiation, and then cover the decoupling laser, which allows for the measurement of the dark counts.
@@
Similarly to the calibration of the A-T splitting regime, the line function with a fixed slope is fitted to the data for both $I_{\text{src}}$ levels. Points lying near the thermal noise and saturation level are omitted to increase the fit accuracy. As both measurements are performed for the same attenuation range, the ratio $I_{2}/I_{1}$ and $I_{2}$ can be determined precisely using fitted intercept line parameters, with their values being $I_{2}/I_{1}=$25.48\pm 0.07\text{\,}$\$ and $I_{2}=$-33.48\pm 0.07\text{\,}\text{dBm}$\$. Thus full response curve spanning from atomic saturation to thermal radiation level is created.
@@
Finally, the scaling factor between the electric field amplitude and measured photon counts is calculated by using fitted intercept line parameters yielding the absolute calibration of the receiver. The measured points, scaled using the aforementioned steps together with the absolute calibration line, the thermal noise, and the comb tooth level are shown in Fig. 2. The colored regions shown in the figure correspond to the measurement regions for both TRA source intensities, with vertical axes representing the number of measured photon counts per second in both cases, and horizontal axes representing the THz attenuation values in both measurements. Based on the absolute calibration, we then find that the electric field amplitude of the measured THz frequency comb tooth is equal to $E_{\text{c}}=$405\pm 8\text{\,}$\$ and the amplitude of thermal noise integrated in the whole conversion bandwidth is equal to $E_{\text{n}}=$36.7\pm 0.7\text{\,}$\$. The calculated values, as well as the entire calibration, are determined without subtracting the dark counts of the detector. It can also be noted based on Fig. 2 that the response to the frequency comb tooth lies in the detector linear response regime, which we define as a linear dependence of the conversion signal on the intensity of the oncoming THz field.
@@ III.2 Tunable broadband readout of frequency comb signal
Next, we visualize the THz frequency comb by scanning the decoupling laser, thus tuning the detector to different teeth (frequency modes) of the comb. During the scanning process both probe and coupling lasers are locked at the optimal frequencies. For the considered transition, we perform a scan over the range of $3\text{\,}$ with a step of $4\text{\,}$ around the atomic resonance. The measurement is performed by detuning the decoupling laser over the frequency range mentioned above and measuring the number of converted photons per second for each of the detunings. As can be seen in Fig. 3 a), the distance between two subsequent peaks is the same as the MaiTai repetition rate $f_{\text{rep}}$ and remains constant across the scan range, with a regime of higher than unity SNR (signal-to-noise ratio) spanning over more than 2 GHz. Additionally, scans without the comb are performed to observe the intensity of converted thermal radiation.
@@
The behavior of the comb frequency scans shown in Fig. 3 can be further understood by comparing them with the numerical predictions. Under the assumption of the linear response, justified by the absolute calibration of the receiver, the expected photon counts as a function of the $\Delta_{1200}$ for the measurement time much longer than $f_{\text{rep}}$ is given as:
@@
where $I_{n}$ is the intensity of each tooth and $\Delta_{\text{THz}}$ corresponds to the frequency distance of the closest tooth from the transition resonance. The conversion gain $\mathcal{G}$ is the ratio between the amplitude of the THz and converted fields [39, 40], in this case proportional to the atomic coherence between $5^{2}\text{D}_{5/2}$ and $5^{2}\text{P}_{3/2}(\text{F}=3)$ states. As the rest of the lasers remain at the constant frequency, we only consider the dependence of $\mathcal{G}$ on the detuning of the THz field and the decoupling laser. Based on the linear assumption, we then fit the theoretical model to the measured comb frequency scan shown in Fig. 3 a) with the thermal photons subtracted. The theoretical model consists of the steady-state solutions for each of the teeth considered, which are then combined according to Eq. (2) to recover the full scan. The measured comb scan with the fitted theoretical predictions can be seen in Fig. 4. It can be noted that the theoretical model and its envelope match the measurements well, proving the applicability of the detector linear response regime. The divergence of the model, visible for the higher decoupling laser detunings, is a result of the influence of other excited states, namely the hyperfine structure of the $5^{2}\text{D}_{5/2}$ state, which has not been taken into account in the simulations.

### v2

@@ Abstract
Terahertz radiation finds an increasing number of applications, yet efficient generation and detection remain a challenge and an active area of research. In particular, the precise detection of weak and narrowband terahertz signals is notoriously difficult. Here, we employ a novel type of single-photon detector based on Rydberg atoms to both detect and calibrate a terahertz frequency comb over an octave-spanning range, yet with a MHz-level selectivity. We calibrate the intensity of the electric field of the comb against the fundamental atomic properties, while achieving the intensity (power) sensitivity down to $45.2\text{\,}$ ($1.84\text{\,}$) within a single mode of the frequency comb, all in a room-temperature operated setup. Our results elucidate the transition of terahertz frequency combs into the quantum regime, enabling high-precision and high-sensitivity spectroscopy. This breakthrough allows terahertz science to better leverage revolutionary techniques developed for optical frequency combs.
@@ II.1 Detection
The first auxiliary method relies upon a 87Rb vapor cell and two optical fields. The laser beams are introduced into the cell from opposing directions and overlaid colinearly. This setup is the typical two-photon excitation scheme utilized in Rydberg atom electrometry, where usually two lasers are used to excite the atoms. The $780\text{\,}$ probe beam tuned to the D2 line is scanned in frequency and, after interaction with atomic vapors is shone upon an avalanche photodiode (Thorlabs APD430A) to obtain the probe transmission spectrum. By the observation of the EIT (electromagnetically induced transparency) effect, introduced by the counter-propagating $483\text{\,}$ coupling laser tuned to the $5^{2}\text{P}_{3/2}(\text{F}=3)\rightarrow 27^{2}\text{D}_{5/2}$ transition, the setup is capable of performing RF electric field measurements by analyzing the Autler-Townes splitting in the domain of the probe laser detuning [35].
Figure 1: a) Simplified scheme of the setup used for the calibration and detection of THz frequency comb signal. $\mathsf{TRA}$ – automotive radar chip in an RF absorbing box, $\mathsf{HWP}$ – half-waveplate, $\mathsf{QWP}$ – quarter-waveplate, $\mathsf{PD}$ – photodiode, $\mathsf{PM}$ – parabolic mirror, $\mathsf{PCA}$ – photoconductive antenna, $\mathsf{POL}$ – polarizer, $\mathsf{DM}$ – dichroic mirror, $\mathsf{SF}$ – spectral filter. The removable metallic mirror allows switching between signal sources. The TRA beam has slanted polarizers to discard reflections, while the HWP + POL setup has two copies to increase the achievable attenuation. b) State-level diagram of fields and state transitions facilitating the conversion of frequencies centered around various detected frequencies to the 776 nm output optical signal. The targeted THz signal frequencies are $0.125\text{\,}$, $0.160\text{\,}$, $0.182\text{\,}$ and $0.242\text{\,}$. Note that only one THz frequency is targeted at a time – the coupling and decoupling lasers are coarsely detuned to address the relevant transition. Furthermore, by fine-detuning the decoupling laser by $\Delta_{1200}$ it is possible to tune to the frequencies near the chosen transition.
This measurement scheme has a fundamental limitation on the minimal resolvable field, resulting from the finite EIT linewidth, caused by the movement of atoms at room temperature and the arising Doppler broadening. With the objective of measuring much weaker fields, the A-T setup is then extended by introducing a third laser beam, namely the 1278 nm decoupling field tuned to the $26^{2}\text{F}_{7/2}\rightarrow 5^{2}\text{D}_{5/2}(\text{F}=4)$ transition, realizing the full setup, visible in Fig. 1 a).
@@
Such a detection scheme achieves superior detection sensitivity to the intensity of THz fields, down to the thermal noise limit, with the potential for time-resolved measurements and synchronization with a pulsed source. Thanks to having a strong reference field source, it is possible to calibrate the photon count readout to the electric field amplitude measured from A-T splitting. What is more, as is the case with A-T splitting, we are not limited to resonant RF detection and addressing different frequencies is possible by detuning the coupling or decoupling fields. For simplicity, this work explores tuning just the decoupling laser for mapping the bandwidth of a particular RF transition.
@@ II.2 Signal generation
The THz field is introduced from a high-power source, namely an automotive radar chip TRA_120_045 from Indie Semiconductor. While A-T measurements may be performed with the signal entering the vapor cell from the side, the parametric conversion necessitates phase-matching, so the THz field is combined colinearly with optical fields and focused using a parabolic mirror with a through hole for optical beams. This enables colinear propagation of all fields in the vapor cell, and the setup is further improved using custom-made metamaterials, which allow for beam collimation, attenuation, and polarization control [33]. Both lenses and waveplates are made using an FDM (fused deposition modeling) 3D printer and HIPS (high impact polystyrene) filament, while the polarizers are custom-made flex PCBs (printed circuit boards). All of the designs are fine-tuned for the operating transition frequency using FDTD (finite difference time domain) simulations. As indicated in the Fig. 1, the TRA source is enclosed in an RF absorbing box with a small aperture facing the first lens. The enclosure pre-attenuates the source, minimizes reflections from it and ensures that only the radiation passing through the aperture couples to the atoms.
As such, a sufficiently powerful source, like the aforementioned radar chip, is employed for calibration and connecting the results obtained in both detection methods. Subsequently, we can switch to conversion detection and study lower intensity field sources, in particular, a THz frequency comb. We therefore introduce a PCA (photoconductive antenna – BATOP iPCA-21-05-300-800-h), which, given a bias voltage, emits THz pulses following the intensity characteristic of a femtosecond Spectra-Physics MaiTai laser, which has a repetition rate $f_{\text{rep}}$ of $79.49\text{\,}$. The PCA has a hyperhemispherical silicon lens mounted on top, which partially focuses the output. It is then collimated and sent into the setup with the same path, and a metal sheet is used as a mirror to toggle between the high-power chip and the locally weaker PCA.
@@ III.1.1 A-T splitting calibration
We begin the calibration by measuring the A-T splitting in the probe transmission spectrum. For this measurement, the decoupling laser is covered, and only the effects from the coupling laser and THz field can be seen. The probe transmission spectrum is then measured for attenuations of the THz field between 0 and $1.5\text{\,}$ introduced by two attenuators, each consisting of a half-waveplate and a polarizer. The A-T splitting, created by the THz field, can then be directly observed in the probe transmission spectra as the splitting of the EIT caused by the $483\text{\,}$ laser. Next, we determine the Rabi frequencies $\Omega$ for the corresponding splitting by fitting the numerical solutions of the steady state to the measured probe transmission spectra [37]. Subsequently, knowing the dipole moment $d$ of the considered THz transition, the strength of the incoming field can be expressed in the amplitude of the electric field $E$ by the relation
$E=\frac{h}{|d|}\Omega,$ (1)
where $h$ is Planck’s constant. In the experiment, the dipole moment of the main considered atomic transition, $27^{2}\text{D}_{5/2}\rightarrow 26^{2}\text{F}_{7/2}$, is equal to $d=594a_{0}e$ [38], where $a_{0}$ is the Bohr radius and $e$ is the value of the elemental charge. The calibration of the A-T splitting regime is then determined by fitting a linear function with a fixed slope, equal to 1-1, to the calculated electric field amplitudes as a function of THz field attenuation.
Figure 2: Calibration curve derived for the $27^{2}\text{D}_{5/2}\rightarrow 26^{2}\text{F}_{7/2}$ THz transition from the single photons and A-T splitting measurements, together with the fitted calibration line. and both the comb tooth signal and the thermal noise levels. Colored areas correspond to the measurement series for the TRA source powers $I_{1}$ (green, yellow) and $I_{2}$ (blue). Different plot markers represent A-T splitting (square) and photon counts measurement (circle). Each measurement series involved continuously attenuating the same signal using the same set of HWPs and polarizers. For the attenuation values represented on the horizontal axis, the photon counts from the SPAPD were registered and are displayed on the right vertical axis, while the A-T splitting measurements yielded the electric field values visible on the left one. The strong TRA mode measurements had to be performed with a neutral density (ND) filter put before the photon counter, to avoid the regime where detector dead times influence the results. The higher THz powers deviate from the calibration line due to the atomic saturation of the conversion process, while the much slower deviations in the high attenuation regime are caused by the imperfections of the used HWPs and polarizers. For reference, the dashed lines show the measured photon counts of the converted frequency comb signal and thermal radiation.
@@ III.1.2 Photon counts calibration
To perform the measurements of converted photon counts, the probe laser is locked at the transition resonant frequency, the decoupling laser is turned on, and both the coupling and decoupling lasers are tuned to maximize the conversion signal coming from the comb tooth closest to the resonance. Then, the photon count measurements of the converted signal are performed for the attenuation range spanning from 0 to 32 dB via the setup of HWPs and polarizers. To further extend the calibration range, the measurements are done for two TRA source power levels, which we will label as: $I_{1}$ used for the A-T splitting measurements, and $I_{2}$, for which the A-T splitting cannot be measured in the probe transmission spectrum. The $I_{1}$ power according to the manufacturer’s documentation is equal to $I_{1}=$-7\text{\,}\text{dBm}$\$. However, the real power entering the setup is significantly lower due to the pre-attenuation caused by the small aperture in the enclosure. As the conversion signal due to the $I_{1}$ power saturates the SPAPD, for those measurement series, an additional neutral density (ND) filter is added before the photon counter. Finally, we turn off the TRA chip and measure the photon counts corresponding to the frequency comb tooth closest to the atomic resonance. Additionally, we cover both THz sources and measure the conversion signal corresponding to the thermal radiation, and then cover the decoupling laser, which allows for the measurement of the dark counts.
@@
Similarly to the calibration of the A-T splitting regime, the line function with a fixed slope is fitted to the data for both source power levels. Points lying near the thermal noise and saturation level, defined as a point where the response of the receiver becomes nonlinear, are omitted to increase the fit accuracy. Hence, the linear function is fitted only to points lying between $11.5\text{\,}$ (saturation level) and $45\text{\,}$ (near the thermal level) THz signal attenuation. Both of those threshold points have been determined by finding the point for which additional omission of data points does not impact the fit accuracy. As both measurements are performed for the same attenuation range, the ratio $I_{2}/I_{1}$ can be determined precisely using fitted intercept line parameters, with their values being $I_{2}/I_{1}=$25.48\pm 0.07\text{\,}$\$ which gives $I_{2}=$-32.48\pm 0.07\text{\,}\text{dBm}$\$. This way, a full response curve spanning from atomic saturation to thermal radiation level is created.
@@
Finally, the scaling factor between the electric field amplitude and measured photon counts is calculated by using the fitted intercept line parameters yielding the absolute calibration of the receiver. The measured points, scaled using the aforementioned steps together with the absolute calibration line, the thermal noise, and the comb tooth level, are shown in Fig. 2, with the square and circle plot markers representing the A-T splitting and photon counts measurement accordingly. The colored regions shown in the figure correspond to the measurement regions for both TRA source powers, with vertical axes representing the number of measured photon counts per second in both cases, and horizontal axes representing the THz attenuation values in both measurements. Additionally, the noticeably small amount of measured data points for the A-T splitting measurement can be attributed to the comparably low sensitivity of this measurement technique. Based on the absolute calibration, we then find that the electric field amplitude of the measured THz frequency comb tooth is equal to $E_{\text{c}}=$405\pm 8\text{\,}$\$ and the amplitude of thermal noise integrated in the whole conversion bandwidth is equal to $E_{\text{noise}}=$36.7\pm 0.7\text{\,}$\$. The calculated values, as well as the entire calibration, are determined without subtracting the dark counts of the detector. It can also be noted based on Fig. 2 that the response to the frequency comb tooth lies in the detector linear response regime, which we define as a linear dependence of the conversion signal on the power of the oncoming THz field. Finally, to extend the context of the measured values, we may derive rough estimates of the total power of the measured fields. Assuming that the THz field ($\lambda=$2.38\text{\,}$\$) is focused down to a diffraction-limited area $A=\pi(1.22\lambda N)^{2}\approx 27\ \mathrm{mm}^{2}$ by the parabolic mirror with f-number $N=1$ we get a lower-bounded power for the pre-attenuated TRA (corresponding to $E_{\mathrm{TRA}}=$5\text{\times}{10}^{4}\text{\,}$\$) $P_{\mathrm{TRA}}\approx-30\ \mathrm{dBm}$ and for a single tooth of the comb $P_{c}\approx-72.3\ \mathrm{dBm}$.
@@ III.2 Tunable broadband readout of frequency comb signal
Next, we visualize the THz frequency comb by scanning the decoupling laser, thus tuning the detector to different teeth (frequency modes) of the comb. During the scanning process, both probe and coupling lasers are locked at the optimal frequencies. For the considered transition, we perform a scan over the range of $3\text{\,}$ with a step of $4\text{\,}$ around the atomic resonance. The measurement is performed by detuning the decoupling laser over the frequency range mentioned above and measuring the number of converted photons per second for each of the detunings. As can be seen in Fig. 3 a), the distance between two subsequent peaks is the same as the MaiTai repetition rate $f_{\text{rep}}$ and remains constant across the scan range, with a regime of higher than unity SNR (signal-to-noise ratio) spanning over more than 2 GHz. Additionally, scans without the comb are performed to observe the intensity of converted thermal radiation.
@@
The behavior of the comb frequency scans shown in Fig. 3 can be further understood by comparing them with the numerical predictions. Under the assumption of the linear response, justified by the absolute calibration of the receiver, the expected photon counts as a function of the decoupling laser detuning $\Delta_{1200}$, for the measurement time much longer than $f_{\text{rep}}$ is given as:
@@
where $I_{n}$ is the intensity of each tooth and $\Delta_{\text{THz}}$ corresponds to the frequency distance of the closest tooth from the transition resonance. The conversion gain $\mathcal{G}$ is the ratio between the amplitude of the THz and converted fields [39, 40], in this case proportional to the atomic coherence between $5^{2}\text{D}_{5/2}$ and $5^{2}\text{P}_{3/2}(\text{F}=3)$ states. As the rest of the lasers remain at the constant frequency, we only consider the dependence of $\mathcal{G}$ on the detuning of the THz field and the decoupling laser. Based on the linear assumption, we then fit the theoretical model to the measured comb frequency scan shown in Fig. 3 a) with the thermal photons subtracted. The theoretical model consists of the steady-state solutions for each of the teeth considered, which are then combined according to Eq. (2) to recover the full scan. The measured comb scan with the fitted theoretical predictions can be seen in Fig. 4. It can be noted that the theoretical model and its envelope match the measurements well, proving the applicability of the detector’s linear response regime. The divergence of the model, visible for the higher decoupling laser detunings, is a result of the influence of other excited states, namely the hyperfine structure of the $5^{2}\text{D}_{5/2}$ state, which has not been taken into account in the simulations.